# Terrain Diffusion: A Diffusion-Based Successor to Perlin Noise in Infinite, Real-Time Terrain Generation

### v1

@@ 6.2. Real Latency: Time to First and Second Tile
Because generation of any fixed-size region has bounded cost, generating a contiguous $512\times 512n$ strip is $O(n)$. But neighboring tiles reuse cached context, so the cost for querying the first $512\times 512$ region is larger than for subsequent regions.
@@
Motivated by these facts, we measure latency as the time to first tile (TTFT) and time to second tile (TTST) across resolutions. TTFT denotes the delay from model initialization to the first $\times512$ tile becoming available, reflecting initial setup cost. TTST measures the time to generate an adjacent $512\times 512$ tile thereafter, reflecting interactive exploration performance. In practice, while both metrics are bounded, they vary with the specific region location because the number of intersecting windows differs across positions. To account for this, as well as minor variations between runs, we perform each evaluation 100 times at random locations and report the average and standard deviation.
@@
Metric Seconds
TTFT 7.60 ± 1.00
TTST 2.40 ± 1.06
An F-35, one of the fastest conventional aircraft in service at roughly 550 m/s, would traverse a 512×512 tile at 90 m resolution in about 84 seconds. In that time, Terrain Diffusion can produce 35 additional tiles, which places the latency well within the requirements of real-time streaming for any slower application. Even if the terrain is downsampled substantially, reducing traversal time by more than an order of magnitude, the query rate still remains well within the model’s capacity.
@@ 6.3. Qualitative Analysis
Figure 5 shows 20 $\times1024$ tiles from Terrain Diffusion, all from the same seed used for Fig.  1. The model produces sharp ridges, coherent river basins, smooth transitions, and varied landscapes. No visible tiling artifacts confirm the effectiveness of InfiniteDiffusion.
@@ 7. Discussion
Coherence, hydrology, and diversity.
Terrain Diffusion generates only elevation plus a few coarse climate variables, yet the resulting terrain is highly coherent. Landscapes aligns with climate, valleys reflect plausible erosion patterns, with hydrological consistency that far exceeds procedural noise, although isolated basins still occur. Qualitatively, the model synthesizes a wide range of landforms, from mountain ranges and volcanoes to canyons, fjords, and coastal plains, all without class conditioning. This diversity, combined with seamless infinite generation and real-time inference, shows that diffusion models can capture structured planetary complexity while retaining many or all the important properties of procedural noise.
@@
The main limitation of our method is that each model relies on conditioning from the level above it, and global coherence deteriorates if this sequence is broken. Some top level must therefore be specified externally. In this work, simple Perlin noise is sufficient because continental scale structure is extremely coarse and contains only broad gradients, so Perlin provides a reasonable starting point while remaining seed consistent and controllable. In domains where a procedural prior is not available, the appropriate replacement is less obvious. A learned generator such as InfinityGAN, or any padding-free GAN, may provide a viable solution, since seed consistency is preserved and the simplicity of the distribution at very coarse scales makes the usual drawbacks of GANs much less significant. We experimented with an end-to-end hierarchy using a small padding-free GAN as the base model, but its outputs were largely comparable to Perlin noise and offered reduced controllability. This direction remains open for future work.
Why InfiniteDiffusion works at low $T$.
While InfiniteDiffusion is practically limited to small $T$, we find that the method retains striking coherence even when generation is aggressively truncated to $T=1$ or $2$. We hypothesize that because global structure is largely determined from hierarchical conditioning, overlapping windows only need to enforce local consistency that is already largely aligned through shared noise. Additionally, because the diffusion models are trained on random crops, they learn an approximately translation invariant representation that bases predictions on local patterns, which are shared across overlapping windows, rather than absolute position, further reducing seams. Together, these factors make the method robust even under minimal iterative refinement.
@@ 8. Future Work
Adding features to the hierarchy is a natural next step. The coarse model, the base model, or both could incorporate additional variables such as soil properties, other climatic variables, or satellite imagery, enhancing control and enabling additional downstream applications. Resolution can also be extended beyond the current hierarchy by adding further refinement stages. While increased resolution does require more work for the same real-world area, traversal speed and viewing distance usually scale with resolution. In this case, the lower throughput required at low resolutions typically balances the extra work at high resolutions, making high-resolution generation highly efficient. Finally, the InfiniteDiffusion formulation itself is not specific to terrain. Any domain that can be decomposed into overlapping tiles can adopt the same sampling strategy, including textures, maps, and large environments in general.
@@ 9. Conclusion
We have presented Terrain Diffusion, a diffusion-based framework for coherent, real-time terrain generation across planetary scales. By reformulating MultiDiffusion for unbounded domains and introducing the Infinite Tensor framework, we enable seed-consistent, random-access synthesis of infinite terrain with constant memory and compute. A hierarchical stack of diffusion and consistency models unifies planetary organization with local realism, producing continuous landscapes far beyond the reach of procedural noise. Together, these components position diffusion models as a practical foundation for procedural worldbuilding.
Appendix A Formal Properties of InfiniteDiffusion
A.1. Preliminaries and Notation
@@ A.2. Seed Consistency
Definition A.1 (Seed-consistent generative process).
@@
Lemma A.2 (InfiniteDiffusion is seed-consistent).
@@
Lemma A.2 shows that InfiniteDiffusion has a well-defined deterministic output for a fixed seed. We now show that the lazy query algorithm is consistent with this definition, and is therefore also seed-consistent.
Lemma A.3 (Algorithm Consistency).
@@
Lemma A.4 (Correctness of a single query).
@@
By Lemma A.3, after the query finishes we have
@@
Theorem A.5 (Seed consistency of the algorithm).
@@
By Lemma A.2, for each $s$, $t$, and finite $R$, $J_{t}^{(s)}[R]$ is uniquely determined by $(s,R)$, so the process defined by (3) is seed-consistent. By Lemma  A.4, the algorithm is equivalent, and therefore also seed-consistent. ∎
Theorem A.5 formalizes the informal argument in the main text: once a seed is fixed, the entire trajectory $(J_{t}^{(s)})_{t=0}^{T}$ is fully determined, and the lazy querying and caching scheme merely memoizes a deterministic computation without affecting its outcome. In particular, querying regions in different orders or repeating a query for the same region cannot change the result.
A.3. Constant-Time Random Access
@@
Lemma A.6 (Recursive cost bound).
@@
Theorem A.7 (Uniform bound on cost for window regions).
@@
For any window index $i$, Theorem A.7 gives a uniform bound $C_{0}(R_{i})\leq K$ on the number of calls to $\Phi$ needed to answer a query for $J_{0}[R_{i}]$, starting from an empty cache. This bound depends only on $T$ and $M$, which are constants, and not on $i$ or any regions previously processed. In standard algorithmic notation, this means that the time complexity of a query $J_{0}[R]$ is $O(1)$, when $R$ is a window region.
@@
Combined with seed consistency, Theorem  A.7 formally justifies the claim that users can jump to arbitrary locations in the infinite world and query any window region efficiently, without needing to generate intermediate tiles and without affecting the content of the regions. When we assume that any region of fixed size has a bounded number of window overlaps, the same argument applies to arbitrary regions of bounded size, not just individual windows.
A.4. Parallelization
@@
Lemma A.8 (Parallel window updates at a fixed timestep).
@@
Theorem A.9 (Parallelization of finite query sets).
@@
By Lemma A.2, once $s$ is fixed the values $J_{u}^{(s)}[R]$ are uniquely determined for all $u$ and all finite regions $R$. In particular, the set of model evaluations that appear in the recursive computation of $\{J_{t}^{(s)}[R^{(k)}]\}_{k=1}^{m}$ is fixed; only their order of execution is not.
@@
Once all these regions are available, the corresponding values $J_{t}^{(s)}[R_{i}]$ for $i\in I_{t}$ are determined and can be formed without further calls to $\Phi$. At this point, Lemma A.8 implies that all remaining evaluations
@@
Theorem A.9 shows that for any fixed seed $s$, timestep $t$ and finite collection of regions $\mathcal{R}$, the recursive computation of $\{J_{t}^{(s)}[R]:R\in\mathcal{R}\}$ admits a bounded-depth parallel schedule of diffusion model evaluations. Combined with seed consistency, this formalizes the claim in the main text that all diffusion model evaluations required to serve any finite batch of queries $J_{t}^{(s)}[R]$ can be performed in parallel, up to the intrinsic sequential dependence across diffusion steps.

### v2

@@ 6.2. Latency: Time to First and Second Tile
Because generation of any fixed-size region has bounded cost, generating a contiguous n-length strip is $O(n)$. But neighboring tiles reuse cached context, so the cost for querying the first region is larger than for subsequent regions.
@@
Motivated by these facts, we measure end-to-end latency as the time to first tile (TTFT) and time to second tile (TTST) across resolutions. TTFT denotes the delay from model initialization to the first tile becoming available, reflecting initial setup cost. TTST measures the time to generate an adjacent tile thereafter, reflecting interactive exploration performance. While both metrics are bounded, they vary with the specific region location because the number of intersecting windows differs across positions. To account for this, we perform 1000 runs at random locations and report the average and standard deviation.
@@
Region Size TTFT TTST
$512\times 512$ $1.72\pm 0.19$ $0.66\pm 0.18$
$256\times 256$ $1.25\pm 0.19$ $0.26\pm 0.18$
[…]
@@ 6.3. Qualitative Analysis
Figure 4 shows 20 $\times1024$ tiles from Terrain Diffusion, all from the same seed used for Fig.  1. The model produces sharp ridges, visually coherent river basins, smooth transitions, and varied landscapes. No visible tiling artifacts confirm the effectiveness of InfiniteDiffusion.
@@ 7. Discussion
The main limitation of our method is that each model relies on conditioning from the level above it, and global coherence deteriorates if this sequence is broken. Some top level must therefore be specified externally. In this work, simple Perlin noise is sufficient because continental-scale structure is extremely coarse, allowing Perlin to provide a reasonable starting point while remaining seed-consistent and controllable. In domains where a procedural prior is unavailable, a learned generator such as InfinityGAN (Lin et al., 2022) may provide a viable solution. Since our hierarchy can refine and upsample a coarse signal by several orders of magnitude, the traditional limitations of GANs are largely mitigated, as they only need to provide the coarsest global gradients. We experimented with an end-to-end hierarchy using a small padding-free GAN as the base model, which preserved the functional properties of our system, but it offered reduced controllability compared to noise-based alternatives. This direction remains open for future work.
Comparison to procedural noise.
Although Terrain Diffusion retains the formal guarantees of procedural noise, it fundamentally diverges in fidelity and computational cost. By leveraging deep generative models, our method captures complex structural and hydrological realism that noise functions approximate only as superficial textures, with the added flexibility of retraining on new datasets to target different environments. While this comes at the cost of raw throughput compared to the microsecond-level latency of noise, our setup time of 1.72 seconds and steady-state generation time of 0.66 seconds remain highly practical. Ultimately, we do not view this as a total replacement; procedural noise retains decisive advantages in speed, interpretability, and zero-data requirements, suggesting that future systems will likely employ a hybrid approach where noise handles simple, low-latency tasks and learned models manage high-fidelity, complex structure.
@@ 8. Future Work
Adding features to the hierarchy is a natural next step. The coarse model, the base model, or both could incorporate additional variables such as soil properties, other climatic variables, or satellite imagery, enhancing control and enabling additional downstream applications. Resolution can also be extended by adding further refinement stages. While increased resolution does require more work for the same real-world area, traversal speed and viewing distance often scale with resolution. In this case, the lower throughput required at low resolutions typically balances the extra work at high resolutions, making high-resolution generation highly efficient. Finally, the InfiniteDiffusion formulation itself is not specific to terrain. Any domain that can be decomposed into overlapping tiles can adopt the same sampling strategy, including textures, maps, and large environments in general.
@@ 9. Conclusion
We have presented Terrain Diffusion, a diffusion-based framework for coherent, real-time terrain generation across planetary scales. By reformulating MultiDiffusion for unbounded domains and introducing the Infinite Tensor framework, we enable seed-consistent, random-access synthesis of infinite worlds with constant memory and compute. A hierarchical stack of diffusion and consistency models unifies planetary organization with local realism, producing continuous landscapes far beyond the reach of procedural noise. Together, these components position diffusion models as a practical foundation for procedural worldbuilding.
Appendix A Extended Results
A.1. Additional Qualitative Samples
In Figure  6, we showcase twenty additional 1024 by 1024 regions from Terrain Diffusion. Samples are uncurated, except for automatically excluding regions with more than 50% ocean in the coarse map. We include additional details on overall elevation range, and climate variables, in the top left of each sample.
A.2. Additional Details on FID Calculations
Below we provide pseudocode for the normalization scheme used to convert arbitrary elevation values to the 0-255 range required for FID calculation.
Algorithm 2 Normalizing heightmaps for FID
Inputs:
[…]
$I_{\min}\leftarrow\min(I)$ $\triangleright$ Per-image minimum
$I_{\max}\leftarrow\max(I)$ $\triangleright$ Per-image maximum
$I_{\text{range}}\leftarrow\max(I_{\max}-I_{\min},255)$ $\triangleright$ Ensure scaling factor $\geq 255$
$I_{\text{mid}}\leftarrow(I_{\min}+I_{\max})/2$
$I_{\text{norm}}\leftarrow\text{clamp}\left(\left(\frac{I-I_{\text{mid}}}{I_{\text{range}}}+0.5\right)\times 255,0,255\right)$
$O\leftarrow\text{repeat}(I_{\text{norm}},\text{channels}=3)$
Output: $O$ cast to uint8
For Perlin blending (Jain et al., 2023), our algorithm is not identical to the one used in the original source, since the original source targets a different, zero-centered dataset. For fair comparison, we adapt the method for our dataset. While the original algorithm applies Perlin noise with a static distribution, we utilize an adaptive distribution that is centered around the mean of the surrounding elevation tiles, and scaled by the standard deviation of these tiles. We utilize linear interpolation to smoothly interpolate between the distributions of neighboring tiles.
A.3. Additional Performance Metrics
In addition to the latency metrics provided in the main text, we also report peak VRAM usage and additional latency metrics for varying region sizes. Our results are in Table  4. For all performance calculations, we initially compile the models with torch.compile(), and use full-precision inference (fp32), since it provided the best performance. VRAM usage reduces to 1846 MB with half-precision floats. We measure latency end-to-end as observed through the same API exposed to applications, including all I/O and system overhead.
Appendix B Formal Properties of InfiniteDiffusion
B.1. Preliminaries and Notation
@@ B.2. Seed Consistency
Definition B.1 (Seed-consistent generative process).
@@
Lemma B.2 (InfiniteDiffusion is seed-consistent).
@@
Lemma B.2 shows that InfiniteDiffusion has a well-defined deterministic output for a fixed seed. We now show that the lazy query algorithm is consistent with this definition, and is therefore also seed-consistent.
Lemma B.3 (Algorithm Consistency).
@@
Lemma B.4 (Correctness of a single query).
@@
By Lemma B.3, after the query finishes we have
@@
Theorem B.5 (Seed consistency of the algorithm).
@@
By Lemma B.2, for each $s$, $t$, and finite $R$, $J_{t}^{(s)}[R]$ is uniquely determined by $(s,R)$, so the process defined by (3) is seed-consistent. By Lemma  B.4, the algorithm is equivalent, and therefore also seed-consistent. ∎
Theorem B.5 formalizes the informal argument in the main text: once a seed is fixed, the entire trajectory $(J_{t}^{(s)})_{t=0}^{T}$ is fully determined, and the lazy querying and caching scheme merely memoizes a deterministic computation without affecting its outcome. In particular, querying regions in different orders or repeating a query for the same region cannot change the result.
B.3. Constant-Time Random Access
@@
Lemma B.6 (Recursive cost bound).
@@
Theorem B.7 (Uniform bound on cost for window regions).
@@
For any window index $i$, Theorem B.7 gives a uniform bound $C_{0}(R_{i})\leq K$ on the number of calls to $\Phi$ needed to answer a query for $J_{0}[R_{i}]$, starting from an empty cache. This bound depends only on $T$ and $M$, which are constants, and not on $i$ or any regions previously processed. In standard algorithmic notation, this means that the time complexity of a query $J_{0}[R]$ is $O(1)$, when $R$ is a window region.
@@
Combined with seed consistency, Theorem  B.7 formally justifies the claim that users can jump to arbitrary locations in the infinite world and query any window region efficiently, without needing to generate intermediate tiles and without affecting the content of the regions. When we assume that any region of fixed size has a bounded number of window overlaps, the same argument applies to arbitrary regions of bounded size, not just individual windows.
B.4. Parallelization
@@
Lemma B.8 (Parallel window updates at a fixed timestep).
@@
Theorem B.9 (Parallelization of finite query sets).
@@
By Lemma B.2, once $s$ is fixed the values $J_{u}^{(s)}[R]$ are uniquely determined for all $u$ and all finite regions $R$. In particular, the set of model evaluations that appear in the recursive computation of $\{J_{t}^{(s)}[R^{(k)}]\}_{k=1}^{m}$ is fixed; only their order of execution is not.
@@
Once all these regions are available, the corresponding values $J_{t}^{(s)}[R_{i}]$ for $i\in I_{t}$ are determined and can be formed without further calls to $\Phi$. At this point, Lemma B.8 implies that all remaining evaluations
@@
Theorem B.9 shows that for any fixed seed $s$, timestep $t$ and finite collection of regions $\mathcal{R}$, the recursive computation of $\{J_{t}^{(s)}[R]:R\in\mathcal{R}\}$ admits a bounded-depth parallel schedule of diffusion model evaluations. Combined with seed consistency, this formalizes the claim in the main text that all diffusion model evaluations required to serve any finite batch of queries $J_{t}^{(s)}[R]$ can be performed in parallel, up to the intrinsic sequential dependence across diffusion steps.
Appendix C The Infinite Tensor Framework
To support the computational requirements of InfiniteDiffusion, specifically the need for bounded-memory operations on unbounded domains, we developed infinite-tensor. This open-source Python library [^2] Available at https://github.com/xandergos/infinite-tensor abstracts the management of sliding windows, caching, and dependency chaining, allowing the implementation of diffusion models to remain focused on mathematical operations rather than memory management.
C.1. Core Abstractions
The framework treats an infinite tensor $\mathcal{T}$ not as a stored array of data, but as a lazily evaluated, immutable object defined by a deterministic generator function $f$. For context, we typically store the accumulators $A_{t}$ and $B_{t}$ from Section  3.3 as one infinite tensor with an extra channel for the weights $B_{t}$.
  • Infinite Tensors: A tensor is defined by a shape $(S_{1},\dots,S_{d})$, where any dimension $S_{i}$ may be infinite (represented as None). The tensor is backed by a generator function $f$ rather than raw memory.
  • Windows: $f$ operates on fixed-size windows. A window is defined by its size $(H,W)$, stride $s\in\mathbb{Z}^{2}$ and offset $o\in\mathbb{Z}^{2}$.
  • The TileStore: All tensors are managed by a TileStore, which acts as the central registry for memory management and caching. It ensures that repeated access to the same spatial region retrieves cached data rather than re-triggering computation, and ensures that dependencies between tensors are tracked. TileStore’s may be linked to a file for persistent storage.
C.2. Caching Strategies
The framework currently implements two caching mechanisms for storing an infinite tensor $\mathcal{T}$. These are configured via the cache_method parameter during tensor creation.
Direct Caching (cache_method=’direct’)
This method uses direct memoization of the function $f$. When a window is computed, its output is stored in an LRU (Least Recently Used) cache. Essentially, we store the $x$ in $\mathcal{T}[R_{i}]\leftarrow\mathcal{T}[R_{i}]+x$ and reconstruct $\mathcal{T}[R]$ on demand (assuming $\mathcal{T}=\mathbf{0}$ initially) by revisiting all the window regions $R_{i}$ intersecting $R$. The primary benefit of this method is that users can specify a strict memory limit (e.g., cache_limit=10MB), which is managed through an LRU cache or a similar method. Direct caching is ideal for transient inference where minimal RAM usage is a priority, and persistent storage is undesirable.
Indirect Caching (cache_method=’indirect’)
This method implements the accumulation buffer strategy directly. The outputs of overlapping windows are accumulated into fixed-size storage tiles. Overall disk/memory usage is lower since overlapping regions use shared storage. However, fine-grained eviction is difficult, so disk use is unbounded. Storage is, however, highly efficient, making this method ideal for workflows utilizing persistent storage.
C.3. Dependency Chaining and DAGs
The framework supports the construction of directed acyclic graphs (DAGs) of infinite tensors. A tensor $\mathcal{T}_{child}$ can declare a dependency on $\mathcal{T}_{parent}$. When a region of $\mathcal{T}_{child}$ is requested: 1. The framework calculates the required covering region of $\mathcal{T}_{parent}$. 2. The parent region is fetched (triggering upstream generation if necessary). 3. The parent data is sliced and passed to the child’s generator function $f$.
This allows for the construction of complex hierarchical pipelines (e.g., Coarse Model $\rightarrow$ Latent Model $\rightarrow$ Decoder) where each stage is an infinite tensor depending on the previous one.
Appendix D Dataset Details
For efficient processing, data is downloaded and stored in contiguous $2048\times 2048$ tiles, each covering equal surface area at approximately 90m resolution. To maintain uniform ground resolution, tiles are stretched in longitude so that each pixel represents roughly the same area regardless of latitude. This equal-area tiling ensures consistent scale across training and allows all models to train on data without distortion. We also exclude tiles beyond 60 (absolute) degrees of latitude to focus on higher quality data at mid latitudes. Finally, 80% of the tiles are randomly assigned to the train set, and the remainder are withheld for validation.
All models are trained on random crops drawn from random tiles, with sampling biased so that 99% of tiles contain at least 1% land, as ocean regions are simpler and lower priority. Each crop is randomly flipped and rotated in 90 degree increments to reflect our goal of generating infinite, directionless terrain. We include scripts for reproducing the dataset in our open-source repository: https://github.com/xandergos/terrain-diffusion.
Appendix E Signed Square-Root Transform
The signed square-root transform applies the function
$f(x)=\text{sign}(x)\sqrt{|x|}$
to each pixel of the dataset. After inference, we undo this operation with
$f^{-1}(x)=\text{sign}(x)x^{2}.$
The statistical effects of this transform are visualized in Figure  7.